Lanni Products is a start-up computer software development firm, It curently owns computer equipment worth $25,000 and has cash on hand of $24,000 contributed by Lannis owners. Lanni takes out a bank loan. It receives $36,000 in cash and signs a note promising to pay back the loan over 3 years a-1. Prepare the balance sheet just after it gets the bank loan. Assets Liabilities & Shareholders Equity Bank loan Shareholders equity Cash Computer Total Total a-2. What is the ratio of real assets to total assets? (Round your answer to 2 decimal places.) Ratio of real assets to total assets b-1. Prepare the balance sheet after Lanni spends the $60,000 to develop its software product. Liabilities & Shareholders Equity Bank loar Sharsholders equity Assets Software product Computers Total Total b-2. What is the ratio of real assets to total assets? (Round your answer to 2 decimal places.) Ratio of real assets to total assats
Answer #1

a-1

There was $24000 cash in hand plus the borrowed amount of $36000, cash total becomes $60000

Shareholders equity is Total assets - Total liabilities = 85000 - 36000

a-2

The ratio of real assets to total asset is : Computers / Total assets

=25,000/85,000

=0.2941 or 0.29

b-1

Amount liabilities & shareholders equity Amount Assets Software products 60000 Bank Loan Computers Total 36000 49000 85000 2

All the cash is used to purchase software products of $60000

Rest is all unchanged.

b-2

The ratio of real assets to total asset is : Software products + Computers / Total assets

=85,000/85,000

=1